American Products plans to lay off 10% of its employees in its aerospace division and 15% of its empolooyees in its agricultural division. If altogether 12% of the 3000 employees in these two divisions will be laid off, then how many employees are in each division?
Answer by RAY100(1637)

Let Ae = Aero div, and Ag = Agriculture div
.
.10 Ae + .15 Ag = .12 (3000)
.
and, Ae +Ag +3000,,,,Ae = 3000 -Ag
.
.10 ( 3000 -Ag) + .15 Ag = .12 (3000)
.
300 -.1 Ag +.15 Ag = 360
.
.05 Ag = 60
.
Ag = 1200
.
Ae = 3000 -1200 = 1800
.
checking
.
.10(1800) + .15 ( 1200) = .12 (3000)
.
180 +180 =360,.,,,,,,,ok